OpenStudy (anonymous):

simplify 2/x^2+8x+15 +1/x^2+11x+30

OpenStudy (whpalmer4):

Collect like terms...

zepdrix (zepdrix):

Soccerbabe, you need to make your question a little more clear using brackets. Is this what it's suppose to look like?\[\large \frac{2}{x^2+8x+15}+\frac{1}{x^2+11x+30}\]

OpenStudy (anonymous):

yes that is what it was supposed to look like

zepdrix (zepdrix):

Hmm so it looks like we need to factor the bottom polynomials before we can combine these fractions. Do you know how to break \(x^2+8x+15\) into a pair of binomials? :)

OpenStudy (whpalmer4):

Ah, okay, now you've got to make a common denominator. I would suggest factoring the denominators to make that simpler.

OpenStudy (anonymous):

yes it would be \[(x+5)(x+3)\]

OpenStudy (whpalmer4):

And the other one?

zepdrix (zepdrix):

Good good c: so we have,\[\large \frac{2}{(x+5)(x+3)}+\frac{1}{x^2+11x+30}\]Can you do it with the other one as well? :)

OpenStudy (anonymous):

the other one would be \[(x+5)(x+6)\]?

OpenStudy (whpalmer4):

All right! Someone who doesn't need factoring explained! Let's keep her :-)

zepdrix (zepdrix):

lolol ikr :D

zepdrix (zepdrix):

Ok good, \[\large \frac{2}{(x+5)(x+3)}+\frac{1}{(x+5)(x+6)}\]

OpenStudy (anonymous):

haha thanks !:)

OpenStudy (whpalmer4):

You're in good hands with @zepdrix, I've got to go.

OpenStudy (anonymous):

thank you @whpalmer4 I appreciate it:)

zepdrix (zepdrix):

If you were trying to combine these two fractions,\[\large \frac{1}{3\cdot5}+\frac{1}{3\cdot6}\]See how they both have a \(3\)? We need a 6 in the base of the first fraction, and a 5 in the base of the second. So we would multiply the first by 5/5, and the second by 6/6.\[\large \frac{6}{6}\frac{1}{3\cdot5}+\frac{1}{3\cdot6}\frac{5}{5} \qquad = \qquad \frac{6}{3\cdot5\cdot6}+\frac{5}{3\cdot5\cdot6}\qquad=\qquad \frac{11}{3\cdot5\cdot6}\]Maybe it's a silly example, but this is the idea, we need to do something like this before we can starting adding things on top.

zepdrix (zepdrix):

See in our problem how they both have a factor of \(x+5\) on the bottom?

OpenStudy (anonymous):

actually it was a good example and yes I do:)

zepdrix (zepdrix):

So that's our "\(3\)", it's the value we don't need to worry about. But we need the first fraction to have an \(x+6\) in it's base, and the second to have an \(x+3\) in it's base, so we'll have have to do some multiplication to get these similar bases :)

OpenStudy (anonymous):

so for the x+6 would I need to multiply it by 3 to get it to be 3x+18?

zepdrix (zepdrix):

I shouldn't have said 3, I was just trying to make a connection to the example i gave XD lol

zepdrix (zepdrix):

\[\large \color{royalblue}{\frac{x+6}{x+6}}\cdot\frac{2}{(x+5)(x+3)}+\frac{1}{(x+5)(x+6)}\cdot \color{royalblue}{\frac{x+3}{x+3}}\] See how I'm multiplying the top and bottom by the missing factors?

OpenStudy (anonymous):

yes.

zepdrix (zepdrix):

\[\large \frac{2(x+6)}{(x+6)(x+5)(x+3)}+\frac{1(x+3)}{(x+3)(x+5)(x+6)}\]If this step is confusing, you can let me know. I put the brackets around them to show that we don't want to expand these terms out. We just want to write them as multiplication.

OpenStudy (anonymous):

So then the top of the equation would look like\[3x+15?\]

OpenStudy (anonymous):

after you simplify it anyways:)

zepdrix (zepdrix):

yes very good :) it looks like we have a tiny bit more simplification we can do.

zepdrix (zepdrix):

See how each term on top is divisible by 3? Try factoring that 3 out, and you might notice a similar term on top that you also have on the bottom! :O

OpenStudy (anonymous):

x+5

zepdrix (zepdrix):

yah :O interesting!

zepdrix (zepdrix):

\[\large \frac{3\color{royalblue}{(x+5)}}{\color{royalblue}{(x+5)}(x+6)(x+3)}\]

zepdrix (zepdrix):

So maybe we can do a nice cancellation, yes? :)

OpenStudy (anonymous):

yes, yes we can! :)

zepdrix (zepdrix):

bahhh i gotta go! D: I think you've got this pretty much figured out though yes? c:

OpenStudy (anonymous):

yes I do:) thank you very much for your help! :)

zepdrix (zepdrix):

yay team \c:/